11.2.28 Find the length of the following curve (2t+3)3/2 yet 0sts1 3 2 The length of the curve is (Simplify your answer.)
Answer #1

şdution: (Il-8-28) Given that X = (at+3 ) 3/ x+, yo tu ele, ostsi = t+ , 03021 Here Suppose & = f(t) , y = gut) be the paramesides, we get lley y = t + the Differentiate with respect to 4 on both + torbe ++) = 1+ (et) tence The length of the curve on- lieta)dt = [te + at]. = [(*+261) - 0 = 3 -- [6 = 3 unit *** Thanking you .....